Find the equation for a hyperbola with vertices (2,0) and (-2,0) and asymptote y=1/2x
Answer by lwsshak3(11628)

Find the equation for a hyperbola with vertices (2,0) and (-2,0) and asymptote y=1/2x
**
Standard form of equation for a hyperbola with horizontal transverse axis:
(x-h)^2/a^2-(y-k)^2/b^2=1, (h,k)=(x,y) coordinates of center
For given hyperbola:
center: (0,0)
length of transverse axis=4 (-2 to 2)=2a
a=2
a^2=4
..
Slope of asymptote=1/2=b/a
2b=a
b=a/2=2/2=1
b^2=1
..
Equation:
x^2/4-y^2/1=1